OpenStudy (anonymous):

Evaluate 2xy - m when x = -4, y = 3, and m = 6 please show your work i dont know how to do this and this is worth 10 percent of my grade if i fail i will fail math please help me:(

OpenStudy (whpalmer4):

It is simple matching of numbers and letters, followed by arithmetic. \[x=-4\]\[y=3\]\[m=6\] Now take the equation and replace the letters with the corresponding numbers: \[2xy-m = 2(-4)(3)-(6)\] What does that equal?

OpenStudy (anonymous):

i dont get it can u please show exaple?

OpenStudy (whpalmer4):

I did. All you have to do is the arithmetic. What is 2*(-4)*3 - 6 ?

OpenStudy (anonymous):

-2?

OpenStudy (whpalmer4):

What is 2*(-4)?

OpenStudy (anonymous):

-2?

OpenStudy (whpalmer4):

What is 2*4?

OpenStudy (anonymous):

8?

OpenStudy (whpalmer4):

Yes, 2*4 = 8. 2*(-4) = -8, because a positive number (2) times a negative number (-4) gives you a negative result (-8). A negative number times a negative number gives you a positive result, and of course a positive number times a positive number also gives you a positive result. So, we've gotten the first multiplication done: \[2*(-4)*3 - 6 = (-8)*3 - 6\] What is (-8) * 3?

OpenStudy (anonymous):

-22?

OpenStudy (anonymous):

i mean -24

OpenStudy (whpalmer4):

Yes. What is -24 - 6?

OpenStudy (anonymous):

-28

OpenStudy (whpalmer4):

Try again...

OpenStudy (whpalmer4):

If you're at -24 on the number line, and you go 6 units to the left (subtracting 6 = moving 6 units to the left on the number line), where do you end up?

OpenStudy (anonymous):

can u please just tell me this problem and teach me on the other one cuzz this is very importind and i need to put in befor 3:30:(

OpenStudy (whpalmer4):

-24 - 6 what is 24 + 6? -24 -6 is the same thing, with a - sign in front...

OpenStudy (anonymous):

31

OpenStudy (whpalmer4):

4 + 6 = 10, doesn't it? So 24 + 6 = 20+10 = 30. Similarly, -24 - 6 = -30

OpenStudy (anonymous):

so its -30?

OpenStudy (anonymous):

???

OpenStudy (whpalmer4):

Yes...sorry, didn't notice you had responded.

OpenStudy (anonymous):

now can u please right down the whole problmem in order done i wana see how it fully looks like

OpenStudy (whpalmer4):

\[x=-4\]\[y=3\]\[m=6\]First substitute in value of \(x\): \[2xy-m = 2(-4)y-m\]Now \(y\): \[2xy-m = 2(-4)(3) -m \]Now \(m\): \[2xy-m = 2(-4)(3) - (6)\]Do the arithmetic: \[2xy-m = 2(-4)(3)-(6) = (-8)(3)-6 = -24 - 6 = -30\]

OpenStudy (anonymous):

Before anything else you need to learn basic arithmetic, this is a elementary school skill. You should be able to add numbers under 100 together easily. For example: 58+12 = 70 70 - 30=40 39-11=28 11-12=-1 -2-45=-47 You NEED to be able to do this quickly and accurately. You can't add 24 and 6 and get 31... You need to think what 4+6 is...10! So how could the number end in 1? From there the problem is asking to make a simple substitution. If I have: a*b - c with a = 2, b = -3, and c = 4 then it's simply 2*(-3) - 4 = -6-4=-10. Apply this to your current problem and any similar problem.